# Mitigating Text Toxicity with Counterfactual Generation

### v1

@@ 3.2 Toxicity mitigation with counterfactual generation
Textual counterfactual generation methods can be of two types: Text editing heuristics and Counterfactual generation with large language models Bhan et al. (2023a). Text editing heuristics address textual counterfactual generation by slightly modifying the input text whose prediction is to be explained. Important tokens are targeted and modified with mask language models to switch the outcome of the classifier, making this approach very similar to the M&R way of detoxifying text. Text editing heuristics methods differ in the way they target important tokens and the language model used to modify the initial text. Regarding the former methods, they mostly target tokens to be modified by applying LFI methods to the classifier that has to be explained. For example, CLOSS Fern and Pope (2021) applies Ablation, MiCE Ross et al. (2021) gradient-based approaches, CREST Treviso et al. (2023) leverages rationalization methods, and TIGTEC Bhan et al. (2023a) employs KernelSHAP or Self-attention. Next, mask language models used to perform mask infilling are essentially T5, RoBERTa or BERT. This way, M&R methods and Text editing heuristics differ only in the way they target toxic POS. As mentioned in the previous section, the former detect tokens to change by either the use of prior knowledge (vocabulary) or from a by-design interpretable model (logistic model), or from expert-anti-expert disagreement, whereas Text editing heuristics apply LFI methods to a toxicity classifier. Finally, text is modified using the same language models. We propose to group these two kinds of approaches together under the name of "target-then-replace" methods. Figure 3 shows the whole target-then-replace counterfactual detoxification process, where the detoxification finally consists in performing the two following token changes: f**k heck and b***h girl
@@
Counterfactual generation with large language models (CF-LLM) methods build counterfactual examples by leveraging pre-trained generative language models in the same way as TST text detoxification. These methods differ in the language model used to generate text and the way the model is steered towards a specific objective. For example, CASPer Madaan et al. (2022) learns perturbations to steer text generation with BART Lewis et al. (2020) and Polyjuice fine-tunes GPT-2 to generate counterfactual examples. This way, TST and CF-LLM methods differ mainly in the way generative language models are steered towards a specific style or label. We group these two families of approaches together under the name of "steered text generation" methods. Figure 4 summarizes the connection between TST text detoxification and Counterfactual generation with large language models on the one hand, and M&R methods and Text editing heuristics on the other.
@@
Finally, counterfactual generation and text detoxification can be (1) defined in the same way with the objective to find a small change to reach a target state, (2) categorized in two similar families of methods, namely steered text generation and target-then-replace and (3) evaluated with common metrics to assess accuracy, proximity and plausibility.
@@ 4 Experimental Settings
This section presents the automatic and human experimental studies conducted across three datasets to perform toxicity mitigation through counterfactual generation.
@@ 4.1 Experimental protocol
We perform toxicity mitigation on three toxicity datasets from Hallinan et al. (2023). Microagression.com (MAgr) is a public blog containing socially-biased interactions with offending quotes. Social Bias Frames (SBF) is a corpus of offensive content from various online sources. We use a subset of SBF from the microaggressions subreddit where the texts have been labeled as harmful by annotators. DynaHate is a dataset of hate comments that are difficult to detect for a hate-speech classifier. Toxicity mitigation is run on texts initially classified as toxic in all three cases. MAgr contains 951 toxic texts, SBF 460 and DynaHate 500. The three datasets are available in the Github project in the initial paper111https://github.com/shallinan1/MarcoDetoxification/tree/main/datasets.
| Dataset | Metric | MaRCo | ParaGeDi | CondBERT | CF-Detox |
| Self-attention | |||||
| MAgr | %ACC | 91.1 | 96.1 | 98.9 | 96.8 |
| SCORE | 0.109 | 0.066 | 0.028 | 0.075 | |
| %S | 65.6 | 36.6 | 50.8 | 71.1 | |
| %CP | 65.3 | 81.4 | 70.3 | 88.8 | |
| PPL | 0.71 | 0.63 | 2.81 | 2.15 | |
| SBF | %ACC | 75.0 | 92.0 | 95.7 | 87.4 |
| SCORE | 0.249 | 0.102 | 0.066 | 0.164 | |
| %S | 68.9 | 39.3 | 67.4 | 88.3 | |
| %CP | 74.2 | 78.2 | 74.8 | 91.4 | |
| PPL | 0.86 | 0.59 | 1.65 | 1.21 | |
| DynaHate | %ACC | 80.2 | 90.8 | 94.2 | 87.6 |
| SCORE | 0.213 | 0.128 | 0.075 | 0.171 | |
| %S | 68.1 | 35.4 | 66.4 | 88.3 | |
| %CP | 65.4 | 75.2 | 70.9 | 87.3 | |
| PPL | 0.73 | 0.43 | 1.50 | 1.24 |
Counterfactual generator and competitors
@@
We instantiate the method proposed in the previous section by choosing as counterfactual generation method TIGTEC Bhan et al. (2023a). TIGTEC is a target-then-replace textual counterfactual generator that implements several LFI methods to target important tokens to be changed. TIGTEC iteratively masks and replaces tokens with a BERT mask language model following a tree search policy based on beam search. TIGTEC has shown great performances, leading to a competitive compromise is terms of success rate, sparsity and content preservation as compared to other counterfactual generation methods. We use the following settings to run TIGTEC: we first train a BERT classifier on a toxic task dataset from Kaggle222https://www.kaggle.com/datasets/rounak02/imported-data to learn to distinguish toxic texts to non-toxic ones. The classifier performance after training is 94%. This way, counterfactual text detoxification is performed with a classifier that has been trained on a different dataset from the ones used for evaluation.
@@
Toxicity mitigation is then performed by generating counterfactual examples starting from toxic texts to reach a non-toxic state. In the following, we call CF-Detox our toxicity mitigation method based on counterfactual generation. CF-Detox is run in three different versions, targeting toxic POS with three different LFI methods: KernelSHAP, Self-attention and Integrated gradients. In particular, we aggregate self-attention as in the original TIGTEC paper by averaging the attention coefficients related to the CLS token over the attention heads in the last layer of the BERT classifier. CF-Detox is compared to three state-of-the art text toxicity mitigation methods: MaRCo Hallinan et al. (2023), CondBERT and ParaGeDi Dale et al. (2021). The code used to run the all the methods is not provided for anonymity reasons and will be available upon acceptance.
Automatic evaluation
We use the 5 metrics previously introduced to assess toxicity mitigation.
@@
In particular, the toxicity metrics are based on a pre-trained toxicity classifier. The library used to import the pre-trained toxicity classifier is transformers and the model backbone is toxic-bert. This toxicity classifier is different from the one used to steer counterfactual toxicity mitigation. The success rate is computed with the accuracy (%ACC) from the classifier to binary assess if the evaluated text is toxic or non-toxic. %ACC is defined as the number of non-toxic texts over the total number of evaluated texts, with respect to the pre-trained toxicity classifier. The average toxicity score (SCORE) is obtained from the last layer of the classifier before the softmax layer.
@@
Text plausibility is measured with the perplexity score Jelinek et al. (2005) and compared to the perplexity of the original text (). This way, a PPL score lower than 1 indicates than the text plausibility increases whereas PPL higher than 1 means that the detoxified text is less plausible. This score is computed based on the exponential average cross-entropy loss of Gemma-2B Gemma Team et al. (2024), a recently developed small generative language model outperforming GPT-2 while having approximately the same size. The library used to import the pre-trained model is transformers and the backbone is gemma-2b. Due to the presence of outliers when calculating the entropy used to calculate perplexity, perplexity is aggregated using the median rather than the mean operator.
Human-grounded evaluation
@@ 4.2 Results
Global results
Table 1 shows the obtained experimental results by running each method on the same datasets. In this table, CF-Detox has been run using self-attention as on the initial paper to target toxic POS. For each dataset, CF-Detox leads to the most content preserving texts, with the highest %CP and %S scores. On the other hand, CF-Detox performs in average worse than ParaGeDi and CondBERT in terms of detoxification accuracy and score across all datasets. Still, the toxicity of texts generated by CF-Detox is in average lower than that of MarCo over all text corpora. If ParaGeDi and CondBERT mitigate the most toxicity, the resulting detoxified texts are significantly different from the initial one in terms of sparsity and semantic proximity. CondBERT generates the less plausible text across all datasets and degrade text plausibility whereas ParaGedi and MaRCo improve it. In particular, ParaGeDi produces the most plausible text. This result is linked to the fact that ParaGeDi utilizes a paraphrase language model to generate text that is intended to be plausible. However, ParaGeDi modifies the most the initial text in terms of sparsity.
The high perplexity level of the text generated by CF-Detox can be partially attributed to the mask language model used for generating new text: indeed, it is significantly smaller than the encoder-decoder models used by ParaGeDi and MaRCo for text generation. The model used by CF-Detox is a small 66M parameters DistilBERT for masked language model, whereas each encoder-decoder model used by MaRCo and ParaGeDi to generate text are respectively a 139M parameters BART and a 220M parameters T5. Using a bigger mask language model such as BERT-base or BERT-large would improve the plausibility of the text generated by CF-Detox. CF-Detox still generates significantly more plausible text as compared to CondBERT.
Human evaluation
@@
This way, automatic and human evaluation indicate that TIGTEC applied to a toxicity classifier offers another possible compromise between toxicity, meaning preservation and text plausiblity as compared to other state-of-the-art existing methods.
| Dataset | Metric | CF-Detox | ||
| Self-attention | Integrated gradients | KernelSHAP | ||
| MAgr | %ACC | 96.8 | 97.2 | 97.3 |
| SCORE | 0.075 | 0.071 | 0.071 | |
| %S | 71.1 | 71.1 | 70.4 | |
| %CP | 88.8 | 89.7 | 88.3 | |
| PPL | 2.15 | 2.15 | 2.18 | |
| SBF | %ACC | 87.4 | 86.5 | 87.6 |
| SCORE | 0.164 | 0.162 | 0.156 | |
| %S | 88.3 | 88.6 | 87.5 | |
| %CP | 91.4 | 91.4 | 90.7 | |
| PPL | 1.21 | 1.25 | 1.28 | |
| DynaHate | %ACC | 87.6 | 88.6 | 88.4 |
| SCORE | 0.171 | 0.161 | 0.158 | |
| %S | 88.3 | 88.7 | 87.1 | |
| %CP | 87.3 | 88.4 | 87.6 | |
| PPL | 1.24 | 1.24 | 1.32 | |
Ablation study
@@
Table 2 shows the experimental results obtained by running three different versions of CF-Detox on the three datasets of interest. Each CF-Detox instance is defined by the LFI method used to target toxic POS. Table 2 shows that Self-attention, Integrated gradients and KernelSHAP lead to similar results in terms of toxicity mitigation, content preservation and text plausibility. These results highlight that LFI methods of a different nature (perturbation, attention or gradient-based) can all yield good results.
@@
Toxicity mitigation through counterfactual generation methods like TIGTEC has to be performed by choosing the appropriate LFI method to target toxicity based on the available model information. For example, KernelSHAP is appropriate if no information (internal parameters, gradients) is available about the classifier used to counterfactually mitigate toxicity, due to its model-agnostic nature. On the contrary, if gradients are accessible, the use of Integrated gradients is indicated since it is less computationally costly than KernelSHAP. Finally, if all parameters are accessible, using Self-attention is appropriate because it is available at no cost.
@@
Since TIGTEC gradually masks and replaces the tokens in the original toxic text based on LFI, we postulate that the sparser detoxified texts a LFI method induces, the better its performance, as it targets the most discriminating tokens of the initial text. This way, among the LFI methods accurately mitigating toxicity, Integrated gradients give the most faithful explanations (i.e. target the most accurately toxic POS).
5 On the use and misuse of toxicity detection and mitigation tools

### v2

@@ 3.2 Epistemic Similarities Between Text Detoxification and Counterfactual Generation
Text editing heuristics is a family of CF generation methods that builds CF examples by slightly modifying the input text whose prediction is to be explained. Important tokens are targeted and modified with mask language models to switch the outcome of the classifier, making this approach very similar to the M&R way of detoxifying text. Text editing heuristics methods differ in the way they target important tokens and the language model used to modify the initial text. Regarding the former methods, they mostly target tokens to be modified by applying LFI methods to the classifier that has to be explained. For example, CLOSS Fern and Pope (2021) applies Ablation, MiCE Ross et al. (2021) gradient-based approaches, CREST Treviso et al. (2023) leverages rationalization methods, and TIGTEC Bhan et al. (2023b) employs KernelSHAP or Self-attention. Next, mask language models used to perform mask infilling are essentially T5, RoBERTa or BERT. This way, M&R methods and Text editing heuristics differ only in the way they target toxic POS. As mentioned in the previous section, the former detect tokens to change either by the use of prior knowledge (vocabulary), intrinsically interpretable model (logistic model) or expert-anti-expert disagreement, whereas Text editing heuristics apply LFI methods to a toxicity classifier. Finally, text is modified using the same kind of NLP encoder. We propose to group these two kinds of approaches together under the name of "target-then-replace" methods. Figure 2 shows the raw target-then-replace CF detoxification process, where the detoxification consists in performing the two following token changes: f**k hell and you they, leading to a text certainly detoxified, but not perfectly content preserving.
@@
Counterfactual generation with large language models (CF-LLM) approaches build CF examples by leveraging pre-trained generative language models in the same way as TST text detoxifiers. These methods differ in the language model applied to generate text and the heuristic used to steer the model towards a specific objective. For example, CASPer Madaan et al. (2022) learns perturbations to steer text generation with BART Lewis et al. (2020) and Polyjuice fine-tunes GPT-2 to generate CF examples. Thus, TST and CF-LLM methods differ in the way generative language models are steered towards a specific style or label. We group these two families of approaches together under the name of "steered text generation" methods. Figure 3 summarizes the connection between TST text detoxification and CF-LLM generation on the one hand, and M&R methods and Text editing heuristics on the other.
@@
All in all, CF generation and text detoxification can be (1) defined in the same way with the objective to find a small change to reach a target state, (2) evaluated with common metrics and (3) categorized in two similar families of methods, namely steered text generation and target-then-replace.
3.3 Counterfactual Toxicity Mitigation Methodology
We propose to perform toxicity mitigation with a target-then-replace CF generator with respect to the toxicity classifier. We postulate that, following the notations of Equation 1, toxicity mitigation can be performed by setting as a toxic instance of interest with . Therefore, the objective is to find that minimizes the cost function such that . This way, by denoting as a target-then-replace CF generation method and the LFI method used to target important tokens to be modified, is defined such as . The lack of ground truth detoxified texts is then overcome through the use of as an oracle to guide text detoxification while keeping the non-toxic content. This way, text detoxification through CF generation consists in detecting texts classified as toxic by and generating their related detoxified CF examples.
We also propose to use CFI to refine CF explanations (detoxified texts) and make them sparser, i.e. more content preserving. Since CF generation methods can generate explanations involving unnecessary token modifications (see token “you” in Figure 2), we re-run the CF generator, guided by CFI to optimally target the tokens that have to be modified to perform detoxification. Formally, refining a CF example with a CFI operator consists in generating such as with a CFI operator. We categorize this methodology as a target-then-replace-then-refine approach. The entire Figure 2 traces the whole target-then-replace-then-refine process of CF detoxification.
@@ 4 Experimental Settings
This section presents the experimental studies conducted across three datasets. We compare our approach performing CF toxicity mitigation named CF-Detox to three competitors: MaRCo, CondBERT and ParaGeDi.
@@ 4.1.1 Datasets
We perform toxicity mitigation on three toxicity datasets from Hallinan et al. (2023). Microagression.com (MAgr) is a public blog containing socially-biased interactions with offending quotes. Social Bias Frames (SBF) is a corpus of offensive content from various online sources. We use a subset of SBF where the texts have been labeled as hateful by annotators. DynaHate is a dataset of hate comments that are difficult to detect for a hate-speech classifier. Toxicity mitigation is run on texts initially classified as toxic in all three cases. MAgr contains 425 toxic texts, SBF 557 and DynaHate 662. The three datasets are available in the Github project in the initial paper111https://github.com/shallinan1/MarcoDetoxification/tree/main/datasets.
| Dataset | Metric | CondBert | MaRCo | ParaGeDi | CF-Detox | |
| Bert-kshap | Distilbert-kshap | |||||
| Dynahate | PPL | 1.37 | 0.70 | 0.43 | 1.24 | 1.25 |
| %CP | 68.8 | 68.4 | 71.2 | 86.4 | 86.6 | |
| %S | 64.1 | 67.8 | 31.1 | 86.1 | 86.3 | |
| %ACC | 92.3 | 70.2 | 89.9 | 82.9 | 81.9 | |
| SCORE | 0.10 | 0.30 | 0.14 | 0.21 | 0.23 | |
| MAgr | PPL | 2.59 | 0.72 | 0.69 | 2.04 | 2.16 |
| %CP | 66.2 | 65.2 | 78.4 | 84.6 | 86.1 | |
| %S | 47.3 | 66.0 | 33.0 | 70.1 | 70.1 | |
| %ACC | 98.1 | 82.1 | 91.6 | 95.1 | 91.8 | |
| SCORE | 0.05 | 0.20 | 0.13 | 0.10 | 0.14 | |
| SBF | PPL | 1.48 | 0.82 | 0.59 | 1.27 | 1.26 |
| %CP | 75.6 | 74.7 | 75.6 | 90.2 | 90.4 | |
| %S | 68.1 | 68.1 | 37.3 | 87.6 | 88.4 | |
| %ACC | 95.9 | 70.2 | 91.7 | 87.6 | 84.2 | |
| SCORE | 0.07 | 0.31 | 0.12 | 0.16 | 0.21 | |
4.1.2 Counterfactual generator and competitors
@@
We instantiate our method proposed in the previous section by setting as CF generation method TIGTEC Bhan et al. (2023b). The rationale behind this choice is that TIGTEC achieves the best compromise in terms of accuracy, content preservation and plausibility when compared with other major competitors Bhan et al. (2023b). TIGTEC is a target-then-replace textual CF generator that implements several LFI methods to target important tokens to be changed. It iteratively masks and replaces tokens with a BERT mask language model following a tree search policy based on beam search to minimize a cost function. We use the following settings to run TIGTEC, we first train a neural classifier (BERT or DistilBERT) on a toxic task dataset from Kaggle222https://www.kaggle.com/datasets/rounak02/imported-data to learn to distinguish toxic texts from non-toxic ones. The classifier performance after training is respectively 94.1% and 90.2% for BERT and DistilBERT. This way, CF text detoxification is performed with classifiers that have been trained on a different dataset from the ones used for the evaluation.
@@
Toxicity mitigation is then performed by generating CF examples starting from toxic texts in order to reach a non-toxic state. In the following, we call CF-Detox our new toxicity mitigation method based on CF generation. The cost function driving CF search is defined such as: with where is the probability score of being non-toxic with respect to the trained classifier, is a cosine similarity measure between text embeddings obtained from Sentence Transformers and is a hyperparameter set to 0.3. For each neural model, CF-Detox is run in three different versions, targeting toxic POS with either KernelSHAP (kshap), Self-attention (attention) or Integrated gradients (IG). Self-attention is aggregated as in the original TIGTEC paper by averaging all the attention coefficients related to the CLS token over the attention heads in the last layer of the classifier.
We compute CFI to refine the raw CF examples generated by CF-Detox by explaining the differences between the initial instances and their raw detoxified versions via Integrated gradients. CFI is only applied to CF examples (raw detoxified texts) initially involving at least two token modifications.
4.1.3 Automatic evaluation
@@
We use the 5 metrics previously introduced to assess toxicity mitigation. The toxicity metrics are based on a pre-trained toxicity classifier. The library used to import the pre-trained toxicity classifier is transformers and the model backbone is toxic-bert. This toxicity classifier is different from the one used to steer CF-Detox toxicity mitigation. The success rate is computed with the accuracy (%ACC) from the binary classifier to assess if the evaluated text is toxic or non-toxic. %ACC is defined as the number of non-toxic texts over the total number of evaluated texts, with respect to the pre-trained toxicity classifier. The average toxicity score (SCORE) is obtained from the last layer of the classifier.
@@
Text plausibility is measured with the perplexity score Jelinek et al. (2005) and compared to the perplexity of the original text (). This way, a PPL score lower than 1 indicates than the text plausibility increases whereas PPL greater than 1 means that the detoxified text is less plausible. This score is computed based on the exponential average cross-entropy loss of Gemma-2B Gemma Team et al. (2024), a recently developed small generative language model outperforming GPT-2. The library to import the pre-trained model is transformers and the backbone is gemma-2b. Due to the presence of outliers when calculating the entropy used to calculate perplexity, aggregation is performed with the median rather than the mean operator.
4.1.4 Human-grounded evaluation
@@ 4.2 Results
4.2.1 Global results
| Dataset | Metric | CF-Detox | |||||
| Bert | Distilbert | ||||||
| attention | IG | kshap | attention | IG | kshap | ||
| Dynahate | PPL | 1.34 | 1.26 | 1.25 | 1.43 | 1.24 | 1.25 |
| %CP | 87.4 | 86.3 | 86.4 | 89.2 | 87.0 | 86.5 | |
| %S | 86.8 | 87.4 | 86.1 | 87.4 | 87.4 | 86.3 | |
| %ACC | 81.6 | 84.0 | 82.9 | 79.0 | 81.4 | 81.9 | |
| SCORE | 0.23 | 0.20 | 0.21 | 0.27 | 0.23 | 0.23 | |
| MAgr | PPL | 2.38 | 2.14 | 2.04 | 2.76 | 2.26 | 2.16 |
| %CP | 88.0 | 85.8 | 84.6 | 90.5 | 87.1 | 86.1 | |
| %S | 70.7 | 70.7 | 70.1 | 71.4 | 70.8 | 70.1 | |
| %ACC | 94.4 | 95.5 | 95.1 | 89.0 | 92.0 | 91.8 | |
| SCORE | 0.13 | 0.10 | 0.10 | 0.19 | 0.14 | 0.14 | |
| SBF | PPL | 1.35 | 1.26 | 1.27 | 1.32 | 1.27 | 1.26 |
| %CP | 91.5 | 90.4 | 90.2 | 92.3 | 90.9 | 90.4 | |
| %S | 88.7 | 89.2 | 87.6 | 89.0 | 89.3 | 88.4 | |
| %ACC | 86.4 | 86.5 | 87.6 | 82.4 | 84.0 | 84.2 | |
| SCORE | 0.19 | 0.17 | 0.16 | 0.24 | 0.21 | 0.21 | |
Table 1 shows the experimental results obtained by running each method on the same datasets. In this table, CF-Detox has been run using KernelSHAP to target toxic POS. For each dataset, CF-Detox leads to the most content preserving texts, with the highest %CP and %S scores. On the other hand, CF-Detox performs in average worse than CondBERT in terms of detoxification accuracy and score across all datasets. Still, the toxicity of texts generated by CF-Detox is in average lower than that of MarCo and similar to ParaGeDi over all text corpora. If CondBERT mitigates the most toxicity, the resulting detoxified texts are significantly different from the initial ones in terms of sparsity and semantic proximity. CondBERT generates the less plausible text across all datasets and degrades text plausibility whereas ParaGedi and MaRCo improve it. In particular, ParaGeDi produces the most plausible detoxified text. This result stems from the autoregressive properties of the paraphrase language model used by ParaGeDi to generate text that is intended to be plausible. However, the text generated by ParaGeDi diverges the most from the initial text in terms of sparsity and proximity.
The high perplexity level of the text generated by CF-Detox can be partially attributed to the mask language model used for generating new text. Indeed, it is significantly smaller than the encoder-decoder models used by ParaGeDi and MaRCo for text generation. The model used by CF-Detox is a small 66M parameters DistilBERT for masked language model, whereas each encoder-decoder model used by MaRCo and ParaGeDi to generate text are respectively a 139M parameters BART and a 220M parameters T5. Using a bigger mask language model such as BERT-base or BERT-large would improve the plausibility of the text generated by CF-Detox. CF-Detox still generates significantly more plausible texts as compared to CondBERT.
4.2.2 Human evaluation
@@
This way, automatic and human evaluation underpin the fact that CF-Detox offers another possible compromise between toxicity, meaning preservation and text plausibility as compared to other state-of-the-art existing methods.
4.2.3 Ablation study
@@
Table 2 shows the experimental results obtained by running three different versions of CF-Detox on the three datasets of interest. Each CF-Detox instance is defined by the LFI method used to target toxic POS. Table 2 shows that Self-attention, Integrated gradients and KernelSHAP lead to similar results in terms of toxicity mitigation, content preservation and text plausibility. These results highlight that LFI methods of a different nature (perturbation, attention or gradient-based) can all yield good results, which underpins CF-Detox robustness.
@@
Toxicity mitigation through CF generation methods like TIGTEC has to be performed by choosing the appropriate LFI method to target toxicity based on the available model information. For example, KernelSHAP is appropriate if no information (internal parameters, gradients) is available about the classifier used to counterfactually mitigate toxicity, due to its model-agnostic nature. On the contrary, if gradients are accessible, the use of Integrated gradients is recommended since it is less computationally costly than KernelSHAP. Finally, if all parameters are accessible, using Self-attention is appropriate because it is available at no cost.
@@
Since TIGTEC gradually masks and replaces tokens in the original toxic text based on LFI, we postulate that the sparser detoxified texts a LFI method induces, the better its performance, as it targets the most discriminating tokens of the initial text with respect to . This way, Integrated gradients give the most faithful explanations (i.e. target the most accurately toxicity).
4.2.4 CFI enhancement
| Model | Dataset | Similarity | Sparsity |
| BERT | MAgr | +0.5% | +1.2% |
| SBF | +0.4% | +2.2% | |
| DynaHate | +1.5% | +4.7% | |
| DistilBERT | MAgr | +0.3% | +0.9% |
| SBF | +0.5% | +2.2% | |
| DynaHate | +2.0% | +4.8% |
We refine detoxified texts using CFI following the target-then-replace-then-refine methodology previously introduced. Table 3 shows similarity and sparsity growth rate obtained from CFs obtained by running CF-Detox with KernelSHAP to target toxic POS. On average, CFI increases both similarity and sparsity across all datasets, when CF-Detox is either based on BERT or DistilBERT. However, the gains are higher with the DynaHate dataset (respectively +4.7% and +4.8% uplift in sparsity for BERT and DistilBERT) containing the most toxic texts as compared to MAgr and SBF.
Figure 5 displays the relation between the gain induced by CFI and the number of tokens initially modified in the first version of the CF by focusing on the DynaHate dataset. The number of considered token modifications stops at 5, representing 95% of the CFs initially generated. It turns out that for DistilBERT, the higher the number of initially modified tokens, the higher the sparsity and proximity gain (up to +11% for sparsity). Results are overall more stable for BERT across the number of tokens initially changed.
5 Discussion
In this work we showed that XAI methods can be applied to a toxicity classifier to target toxic POS with LFI and mitigate toxicity with CF generation. CF-Detox enables to find a new compromise in terms of toxicity lowering, content preservation, and textual plausibility. We also showed that CFI could be used to enhance detoxification in terms of proximity and sparsity by following a target-then-replace-then-refine approach. We believe that the latter concerns the XAI community in general, to make CF examples closer to the instance to be explained and sparser.
CF toxicity mitigation is highly dependent on the toxicity classifier used to steer detoxification. If is unreliable and only performs well on its training set, the risk of incorrectly indicating that the text has been detoxified is high. Therefore, the choice of and the toxicity training data must be made with caution to avoid incorrect toxicity assessments during detoxification. The classifier can be fine-tuned on a more specific dataset if the detoxification task is related to a precise kind of toxicity, such as racism or sexism.
CF detoxification has been tested by applying three different versions of TIGTEC with KernelSHAP, Integrated gradients, or Self-attention. There is a wide range of other LFI methods that could be used to target important tokens to explain a toxicity prediction. Besides, other CF generators such as MiCE Ross et al. (2021), CREST Treviso et al. (2023) or CLOSS Fern and Pope (2021) could be used to perform toxicity mitigation. We believe that these CF generation methods could lead to other levels of compromise between toxicity lowering, text plausibility and content preservation.